OpenStudy (anonymous):

stimate how many times larger 4 x 1015 is than 19 x 1012. A) 20 times larger B) 50 times larger C) 100 times larger D) 200 times larger

OpenStudy (anonymous):

the ansewer is b

OpenStudy (cmtboy2):

no answer

OpenStudy (benlindquist):

what is with ur pic

OpenStudy (anonymous):

@DERRICKCURRY Dont just give answer. That is against the code of conduct.

OpenStudy (cmtboy2):

you hot

OpenStudy (anonymous):

no one help me without having a sexy pic

OpenStudy (anonymous):

This is not tinder.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

2x - 3y = 16 3x + 2y = 11 Solve the system of equations. A) (15 1 2 ,5) B) (5,-2) C) (5, 4) D) (-1, -6)

OpenStudy (cmtboy2):

a

OpenStudy (cmtboy2):

medal plz

OpenStudy (anonymous):

do you still need help with this problem?

OpenStudy (anonymous):

Solve this linear system. 2.4 = 0.3 x + 0.4y 5x = 2 + 6y A) x = 4; y = 3 B) x = 3; y = 4 C) x = 4; y = -3 D) x = -4; y = 3

OpenStudy (cmtboy2):

d i think

OpenStudy (anonymous):

I would have to go with D. That is what the google calculator is telling me

OpenStudy (cmtboy2):

a medal

OpenStudy (anonymous):

4.83 x 104 = A) 0.000483 B) 483 C) 4,830 D) 48,300

OpenStudy (cmtboy2):

b

OpenStudy (anonymous):

B

OpenStudy (anonymous):

http://www.usatestprep.com/modules/gallery/files/5/535/535.png What is the solution to system of linear equations graphed here? A) (0,1) B) (1,-1) C) (0,-3) D) (-1,1)

OpenStudy (cmtboy2):

c

OpenStudy (anonymous):

I tried it out on the calculator, but it isnt giving me the right answer.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

http://www.usatestprep.com/modules/gallery/files/5/536/536.png Find the solution to the system of equations graphed here. A) (-1, -3) B) (1, -3) C) (1, 3) D) (3, 1)

OpenStudy (cmtboy2):

a

OpenStudy (anonymous):

Rewrite using the properties of exponents. x-3 A) -x3 B) -3x C) 1 x3 D) 3 x3

OpenStudy (cmtboy2):

d

OpenStudy (anonymous):

Shaun is 4 years older than her sister Charmaine. If the sum of their ages is 16 how old is Charmaine? A) 4 years old B) 6 years old C) 8 years old D) 10 years old

OpenStudy (benlindquist):

b

OpenStudy (cmtboy2):

d

OpenStudy (anonymous):

@cmtboy2 You are just typing random answer. @evoodoochild You shouldnt trust it all. Rewrite using the properties of exponents. \[x^{-3}\] can be written as \[\frac{1}{x^3}\] Which is not d as @cmtboy2 said-